A town has a population of 20000 and grows at 4% every year. What will be the population after 6 years, to the nearest whole number?


1
Expert's answer
2021-02-04T06:32:31-0500

The growth of a population can be regarded as an exponential growth:

"P(n)=P_0 (1+r)^n"

"P(n) -" the population after n years

"P_0 -" initial population

"r -" rate of grow

"n-" time


"P_0=20000"

"r=0.04"

"n=6"


"P(n)=20000*(1+0.04)^6=20000*(1.04)^6=20000*1.265=25306"

Answer: the population after 6 years is 25306.